solve using elimination method
3r-2s=-15
2r+3s=29

Answer by drj(1380)

Solve using elimination method

3r-2s=-15 Equation 1
2r+3s=29 Equation 2

Step 1. Multiply Equation 1 by 3 to both sides of the equation to get Equation 1A below and Equation 2 by 2 to get Equation 2A below to eliminate the s terms when we add the following equations

9r-6s=-45 Equation 1A
4r+6s=58 Equation 2A

Step 2. Adding Equations 1A and 2A will eliminate the s terms.

9r+4r=-45+58
13r=`13
Divide by 13 to both sides of the equation yields r=1

Step 3. Substitute r=1 into Equation 1 to find s

3r-2s=-15 or 3-2s=-15 or -2s=-18. Then s=9.

Step 4. With r=1 and s=9, check Equation 2 if true 2r+3s=29 or 2*1+3*9=29 which is a true statement.

Step 5. ANSWER: The solution is r=1 and s=9.

I hope the above steps were helpful.

For FREE Step-By-Step videos in Introduction to Algebra, please visit http://www.FreedomUniversity.TV/courses/IntroAlgebra and for Trigonometry visit http://www.FreedomUniversity.TV/courses/Trigonometry.

Good luck in your studies!

Respectfully,
Dr J